OpenStudy (anonymous):

Simplify:n squared + 8n + 16 over n squared minus 6n + 9 divided by n + 4 over n minus 3

OpenStudy (anonymous):

\[\frac{n^2+8n+16}{n^2-6n+9}\times \frac{n-3}{n+4}\] is a start, then factor as \[\frac{(n+4)(n+4)(n-3)}{(n-3)(n-3)(n+4)}\] and so some canceling

jhonyy9 (jhonyy9):

(n+4)(n+4) ----------- (n-3)(n-3)

jhonyy9 (jhonyy9):

so this divided by (n+4)/(n-3)

jhonyy9 (jhonyy9):

so will get (n+4)/(n-3)

OpenStudy (anonymous):

Thanks! :)

jhonyy9 (jhonyy9):

yw good luck bye